Test whether the series ∞Σn=0 1/(n^5+x^3) converges uniformly or not

1
Expert's answer
2022-03-14T04:33:50-0400

Alternative method


We consider "{a_n} = \frac{1}{{{n^5} + {x^3}}}\"and "{b_n} = \frac{1}{{{n^5}}}\"     


Here we can see both the series "\sum {a_n}\"and "\sum {b_n}" converge. 


Now


"\begin{array}{l}
\mathop {\lim }\limits_{n \to \infty } \frac{{{a_n}}}{{{b_n}}}\\ = \mathop {\lim }\limits_{n \to \infty } \frac{{\frac{1}{{{n^5} + {x^3}}}}}{{\frac{1}{{{n^5}}}}}\\
 = \mathop {\lim }\limits_{n \to \infty } \frac{{{n^5}}}{{{n^5} + {x^3}}}\\ = \mathop {\lim }\limits_{n \to \infty } \frac{{{n^5}}}{{{n^5}\left( {1 + \frac{{{x^3}}}{{{n^5}}}} \right)}}\\
 = \mathop {\lim }\limits_{n \to \infty } \frac{1}{{1 + \frac{{{x^3}}}{{{n^5}}}}}\\ = \frac{1}{{1 + \frac{{{x^3}}}{{{\infty ^5}}}}} = \frac{1}{{1 + 0}} = 1
\end{array}\"


Therefore, by limit comparison test the series "\sum\limits_{n = 0}^\infty {{a_n}} = \sum\limits_{n = 0}^\infty {\frac{1}{{{n^5} + {x^3}}}} \" converges.